OpenStudy (anonymous):

Express the fraction in lowest terms. 24/54

OpenStudy (zzr0ck3r):

devide top and bottom by 2 and see what ahppens

OpenStudy (zzr0ck3r):

then find a number you can evenly devide the top and bottom of that with

OpenStudy (anonymous):

Factor out both the top and bottom of the fraction: \[\frac{24}{54} = \]

OpenStudy (anonymous):

Try dividing both the numerator and denominator by 6, let us see what you get??

OpenStudy (anonymous):

Lead them to water, don't throw them in lol

OpenStudy (zzr0ck3r):

yes best to find the biggest thing that will devide the top and bottom, but if they are even, I always start with 2 then it can sometimes be easy to see:)

OpenStudy (anonymous):

So its 4/9? does that make it 2/3? I suck at math..

OpenStudy (anonymous):

Why are you taking square root of 4/9??

OpenStudy (zzr0ck3r):

4/9 does not reduce for it to reduce you would have to find a whole number that devides them both.

OpenStudy (zzr0ck3r):

4/6 = 2/3 because 4 devided by 2 is 2 and 6 devided by 2 is 3

OpenStudy (anonymous):

\[\frac{24:6}{54:6}=\frac{4}{9}\]

OpenStudy (anonymous):

Lmfao.. I don't even care. I just wanted to know the answer.. haha

OpenStudy (zzr0ck3r):

dont say such things, getting there is all the fun. this is just new to you. we all suck when its new:)

OpenStudy (anonymous):

This site is here to help you figure out the answers for yourself. If people give you the answers, then you'll learn nothing and you'll fail your tests.

OpenStudy (anonymous):

Hah, well I'm taking my test right now and I don't seem to be failing since you kind people helped. Thanks, bunches.

OpenStudy (zzr0ck3r):

doing and helping are two different things, and if you tell us that your just taking the answers and putting them in your test we might start giving you the wrong answer:P

OpenStudy (theviper):

\[\Huge{\color{red}{\frac{4}{9}}}\]